Two particles move parallel to the x- axis about the orgin with the same amplitude ‘a’ and frequency ‘w’.At a certain instant they are found at a distance ‘a/3’ from the orgin on opposite sides their velocities are in the same direction.what is the phase difference between the two?

Grade:12

1 Answers

Sandeep Bhamoo
5 years ago
Make phesor diagram to silve it for both particle take the angle p,q from hrizontal line you will get sin(p)=sin(q)=1/3 hence phase difference is p+q = 2sin inverse if 1/3. =2sin^(-1){1/3} = sin^(-1){4√2÷9}. Please do approve my answer if correct